OpenStudy (anonymous):

Write the solution in lowest terms. -1/3 + 3/4 =

OpenStudy (anonymous):

\[-\frac{4}{12}+\frac{9}{12}\]

OpenStudy (anonymous):

-4+9=?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

-1/3 + 3/4 = -4/12+9/12 = 5/12

OpenStudy (anonymous):

So what's your answer?

OpenStudy (anonymous):

take the LCM, u get 9-4/12 which is equal to 5/12